# Are Your Models Still Fair? Fairness Attacks on Graph Neural Networks via Node Injections

### v1

@@ Abstract
Despite the remarkable capabilities demonstrated by Graph Neural Networks (GNNs) in graph-related tasks, recent research has revealed the fairness vulnerabilities in GNNs when facing malicious adversarial attacks. However, all existing fairness attacks require manipulating the connectivity between existing nodes, which may be prohibited in reality. To this end, we introduce a Node Injection-based Fairness Attack (NIFA), exploring the vulnerabilities of GNN fairness in such a more realistic setting. In detail, NIFA first designs two insightful principles for node injection operations, namely the uncertainty-maximization principle and homophily-increase principle, and then optimizes injected nodes’ feature matrix to further ensure the effectiveness of fairness attacks. Comprehensive experiments on three real-world datasets consistently demonstrate that NIFA can significantly undermine the fairness of mainstream GNNs, even including fairness-aware GNNs, by injecting merely 1% of nodes. We sincerely hope that our work can stimulate increasing attention from researchers on the vulnerability of GNN fairness, and encourage the development of corresponding defense mechanisms.
**footnotetext: Authors are affiliated with the National Engineering Research Center for Big Data Technology and System, Service Computing Technology and Systems Laboratory, Cluster and Grid Computing Lab, School of Computer Science and Technology, Huazhong University of Science and Technology.footnotetext: Hong Huang is the corresponding author.
@@ 1 Introduction
Due to the strong capability in understanding graph structure, Graph Neural Networks (GNNs) have achieved much progress in graph-related domains such as social recommendation (Song et al., 2023, 2022) and bioinformatics (Coşkun and Koyutürk, 2021; Rozemberczki et al., 2022). Nevertheless, despite the impressive capabilities demonstrated by GNNs, more and more in-depth research has revealed shortcomings in the fairness of GNN models, which greatly restricts their applications in the real world.
@@
In fact, studies (Dai and Wang, 2021; Zhang et al., 2022) have found that the biases and prejudices existed in training data would be further amplified through the message propagation mechanism of GNNs, leading to model predictions being correlated with certain sensitive attributes, such as gender and race. Such correlations are usually undesired and can result in fairness issues and societal harm. For instance, in online recruitment, a recommender based on GNNs may be associated with the gender of applicants, leading to differential treatments towards different demographics and consequently giving rise to group unfairness. To address fairness issues in GNNs, researchers have proposed solutions such as adversarial learning (Dai and Wang, 2021; Singer and Radinsky, 2022), data augmentation (Ling et al., 2023; Luo et al., 2023) and others, which have achieved promising results.
@@
However, recent research in the machine learning domain indicates that fairness is actually susceptible to adversarial attacks (Chhabra et al., 2023; Mehrabi et al., 2021; Solans et al., 2020). Given this, we cannot help but wonder: “Is the fairness of GNN models also highly vulnerable?” For example, in e-commerce, if attackers could exacerbate performance disparities between male and female user groups by attacking GNN-based recommendation models, they could ultimately cause the e-commerce platform to provoke dissatisfaction from specific user demographics and gradually lose its appeal among these users. Several studies (Hussain et al., 2022; Kang et al., 2023; Zhang et al., 2024) have explored the vulnerability of GNN fairness and proposed effective attack strategies. Unlike conventional attacks, these fairness attacks aim to undermine GNN fairness without excessively compromising its utility. However, all these works require altering connectivity between existing nodes, whose authority is typically limited in the real world, such as modifying the relationship between real users. In contrast, injecting fake nodes into the original graph is a more practical way to launch an attack without manipulating the existing graph (Ju et al., 2023; Sun et al., 2020; Tao et al., 2021), which is still under-explored in the field of GNN fairness attack. To address this gap, we aim to be the first to launch an attack on GNN fairness via node injection, examining their vulnerabilities under such a more realistic setting.
@@
Specifically, launching a node injection-based fairness attack on GNNs is non-trivial, whose challenges can be summarized as follows: RQ1: How to determine the node injection strategy? The node injection can be decomposed into two steps, including selecting appropriate target nodes and connecting the injected nodes with them, both will impact the effectiveness of the attack. RQ2: How to determine the features of the injected nodes after node injection? Like the real nodes, injected nodes will also participate in the message propagation process of GNNs, thereby affecting their neighbors and even the whole graph. Given the key role of massage propagation in GNN fairness (Yang et al., 2024; Zhu et al., 2023), proposing suitable strategies to determine the features of injected nodes is also important.
@@
To address these challenges, we propose a gray-box poisoning attack method namely Node Injection-based Fairness Attack (NIFA) during the GNN training phase. In detail, for the two steps in the first challenge, NIFA innovatively designs two corresponding principles. The first is the uncertainty-maximization principle, which asks to select real nodes with the highest model uncertainty as target nodes for injection. The idea is that nodes with higher uncertainty are typically more susceptible to attacks, thereby ensuring the attack’s effectiveness. After selecting target nodes, NIFA follows the second principle, the homophily-increase principle, to connect target nodes with injected nodes. This principle aims to deteriorate GNN fairness by enhancing message propagation within sensitive groups (Luo et al., 2023; Yang et al., 2024). For the second challenge, multiple novel objective functions are proposed after node injection to guide the optimization of the injected nodes’ features, which could further impact the victim GNN’s fairness from a feature perspective. In summary, our contributions are as follows:
@@ 2 Related work
Fairness on GNNs. Researchers have discovered various fairness issues of GNNs, which often lead to societal harms (Luo et al., 2023; Wang et al., 2022a) and performance deterioration (Liu et al., 2023; Tang et al., 2020) in practical applications. Algorithmic fairness on GNNs can be categorized into two main types based on the definition: individual fairness (Agarwal et al., 2021; Dong et al., 2021) and group fairness (Dai and Wang, 2021; Dong et al., 2022; Wang et al., 2022a; Zhu et al., 2024). Individual fairness requires that similar individuals should receive similar treatment, while group fairness aims to protect specific disadvantaged groups (Li et al., 2021). In detail, many researchers have delved into studies focusing on fairness grounded in sensitive attributes. For instance, Dai et al. (Dai and Wang, 2021) reduce the identifiability of sensitive attributes in node embeddings through adversarial training to enhance fairness. FairVGNN (Wang et al., 2022b) goes a step further by introducing a feature masking strategy to address the problem of sensitive information leakage during the feature propagation process in GNNs. Graphair (Ling et al., 2023) achieves fairness through an automated data augmentation method and FairSIN (Yang et al., 2024) designs a novel sensitive information neutralization method for fairness. Beyond fairness related to sensitive attributes, some researchers also direct attention to fairness related to graph structures, like DegFairGNN (Liu et al., 2023) and Ada-GNN (Luo et al., 2022). In this work, we mainly focus on attacks on the group fairness of GNNs based on sensitive attributes.
@@
Attacks on GNNs. Finding out potential vulnerabilities thus improving the security of GNNs remains a pivotal concern in the field of trustworthy GNNs (Zhang et al., 2022). From the perspective of attackers, they aim to compromise the GNNs’ performance on graph data via manipulating graph structures (Ma et al., 2021; Sun et al., 2020; Zou et al., 2021), node attributes (Zügner et al., 2018), or node labels (Liu et al., 2019). Among these methods, node injection attacks, given the attackers’ limited authority to manipulate the connectivity between existing nodes, emerge as one of the most prevalent methods (Ju et al., 2023; Sun et al., 2020; Tao et al., 2021). However, existing attacks mainly focus on undermining GNN’s utility, with little attention to the vulnerability of GNN fairness. FA-GNN (Hussain et al., 2022), FATE (Kang et al., 2023), and G-FairAttack (Zhang et al., 2024) stand out as the few ones that we are aware of to explore attacks on GNN fairness. FA-GNN’s empirical findings suggest that adding edges with certain strategies can significantly compromise GNN fairness without affecting its performance (Hussain et al., 2022). FATE (Kang et al., 2023) formulates the fairness attacks as a bi-level optimization problem and proposes a meta-learning-based framework. G-FairAttack (Zhang et al., 2024) designs a novel surrogate loss with utility constraints to launch the attacks in a non-gradient manner. Nevertheless, all these works require modifying the link structure between existing nodes, which may be prohibited in reality due to the lack of authority.
@@ 3 Preliminary
Notations. A graph is denoted as $\mathcal{G}=(\mathcal{V},\textbf{A},\textbf{X})$, where $\mathcal{V}$ is the node set, and $\textbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ represents the adjacency matrix. $\textbf{X}\in\mathbb{R}^{|\mathcal{V}|\times D}$ denotes the feature matrix, in which $D$ is the feature dimension. Under the settings of node classification, each node $v\in\mathcal{V}$ will be assigned with a label $y_{v}\in\mathcal{Y}$, and a GNN-based mapping function $f_{\theta}:\{\mathcal{V},\mathcal{G}\}\rightarrow\{1,2,...,|\mathcal{Y}|\}^{|\mathcal{V}|}$ with parameters $\theta$ is learned to leverage the graph signals for label prediction, where $\mathcal{Y}$ represents the true label set.
Fairness-related concepts. In alignment with prior works (Dai and Wang, 2021; Dong et al., 2022; Ling et al., 2023; Wang et al., 2022b), we mainly focus on group fairness where each node will be assigned with a binary sensitive attribute $s\in\{0,1\}$, although our attack could also be generalized to the settings of multi-sensitive groups and we leave this as our future work. Based on the sensitive attributes, the nodes can be divided into two non-overlapped groups $\mathcal{V}=\{\mathcal{V}_{0},\mathcal{V}_{1}\}$, and we employ the following two kinds of fairness related definitions:
@@
$\Delta_{SP}=\mathbb{E}|P(\hat{y}=y|s=0)-P(\hat{y}=y|s=1)|$ (3)
$\Delta_{EO}=\mathbb{E}|P(\hat{y}=y|y=y,s=0)-P(\hat{y}=y|y=y,s=1)|$ (4)
Problem definition. In this paper, our goal is to launch fairness-targeted attacks on GNN models through the application of node injection during the training phase, i.e. poisoning attack. Following the line of previous attacks on GNNs (Sun et al., 2020; Zhang et al., 2024), our attack is under the prevalent gray-box setting, where the attackers can obtain the graph $\mathcal{G}$ with node labels $\mathcal{Y}$, and the sensitive information $s$, but can not access the model architecture and parameters $\theta$. Detailed introduction to our attack settings are provided in Appendix B. Specifically, through injecting malicious node set $\mathcal{V}_{I}$ into the graph, the original graph $\mathcal{G}=(\mathcal{V},\textbf{A},\textbf{X})$ is poisoned as $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\textbf{A}^{\prime},\textbf{X}^{\prime})$, where
$\mathcal{V}^{\prime}=\mathcal{V}\cup\mathcal{V}_{I},\enspace\textbf{X}^{\prime}=\left[\begin{array}[]{cc}\textbf{X}\\ \textbf{X}_{I}\\ \end{array}\right],\textbf{X}_{I}\in{\mathbb{R}^{|\mathcal{V}_{I}|\times D}}$ (5)
$\textbf{A}^{\prime}=\left[\begin{array}[]{cc}\textbf{A}&\textbf{V}_{I}\\ \textbf{V}_{I}^{T}&\textbf{A}_{I}\\ \end{array}\right],\textbf{V}_{I}\in{\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}_{I}|}},\textbf{A}_{I}\in{\mathbb{R}^{|\mathcal{V}_{I}|\times|\mathcal{V}_{I}|}}$ (6)
@@
$\begin{array}[]{c@{}l}\mbox{s.t.}&\mathop{\arg\max}\limits_{\theta^{*}}\mathcal{M}(f_{\theta^{*}}(\mathcal{V},\mathcal{G}^{\prime})),\enspace\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\textbf{A}^{\prime},\textbf{X}^{\prime}),\enspace|\mathcal{V}_{I}|\leq b,\enspace deg(v)_{v\in\mathcal{V}_{I}}\leq d\end{array}$
As a poisoning attack, the first constraint in Eq. (7) requires to train the victim model $f_{\theta^{*}}$ with parameters $\theta^{*}$ on the poisoned graph $\mathcal{G}^{\prime}$, so that the predictions of $f_{\theta^{*}}$ are as correct as possible before evaluating the attack performance. The following constraints in Eq. (7) make sure that the proposed attack is unnoticeable and deceptive to the defenders, i.e. the number of injected nodes is below a predefined budget $b$ [^1] Same as the prior work (Sun et al., 2020), we define the perturbation rate as the ratio of injected nodes to the labeled nodes in the original graph, i.e. $|\mathcal{V}_{I}|/|\mathcal{V}_{L}|$, where $\mathcal{V}_{L}$ denotes the labeled node set. and the degrees of injected nodes are constrained by a budget $d$. Our goal is to find a poisoned graph $\mathcal{G}^{\prime}$ to deteriorate the fairness of victim models $f_{\theta^{*}}$ as severely as possible, i.e. maximize the fairness metrics $\Delta_{SP}$ and $\Delta_{EO}$ introduced previously.

### v2

@@ Abstract
Despite the remarkable capabilities demonstrated by Graph Neural Networks (GNNs) in graph-related tasks, recent research has revealed the fairness vulnerabilities in GNNs when facing malicious adversarial attacks. However, all existing fairness attacks require manipulating the connectivity between existing nodes, which may be prohibited in reality. To this end, we introduce a Node Injection-based Fairness Attack (NIFA), exploring the vulnerabilities of GNN fairness in such a more realistic setting. In detail, NIFA first designs two insightful principles for node injection operations, namely the uncertainty-maximization principle and homophily-increase principle, and then optimizes injected nodes’ feature matrix to further ensure the effectiveness of fairness attacks. Comprehensive experiments on three real-world datasets consistently demonstrate that NIFA can significantly undermine the fairness of mainstream GNNs, even including fairness-aware GNNs, by injecting merely 1% of nodes. We sincerely hope that our work can stimulate increasing attention from researchers on the vulnerability of GNN fairness, and encourage the development of corresponding defense mechanisms. Our code and data are released at: https://github.com/CGCL-codes/NIFA.
footnotetext: Hong Huang is the corresponding author. Zihan Luo, Hong Huang, Yongkang Zhou, Jiping Zhang and Hai Jin are affiliated with the National Engineering Research Center for Big Data Technology and System, Service Computing Technology and Systems Laboratory, Cluster and Grid Computing Lab, School of Computer Science and Technology, Huazhong University of Science and Technology.
@@ 1 Introduction
Due to the strong capability in understanding graph structure, Graph Neural Networks (GNNs) have achieved much progress in graph-related domains such as social recommendation [30, 31] and bioinformatics [4, 27]. Nevertheless, despite the impressive capabilities demonstrated by GNNs, more and more in-depth research has revealed shortcomings in the fairness of GNN models, which greatly restricts their applications in the real world.
@@
In fact, studies [5, 42] have found that the biases and prejudices existed in training data would be further amplified through the message propagation mechanism of GNNs, leading to model predictions being correlated with certain sensitive attributes, such as gender and race. Such correlations are usually undesired and can result in fairness issues and societal harm. For instance, in online recruitment, a recommender based on GNNs may be associated with the gender of applicants, leading to differential treatments towards different demographics and consequently giving rise to group unfairness. To address fairness issues in GNNs, researchers have proposed solutions such as adversarial learning [5, 28], data augmentation [17, 22] and others, which have achieved promising results.
@@
However, recent research in the machine learning domain indicates that fairness is actually susceptible to adversarial attacks [3, 26, 29]. Given this, we cannot help but wonder: “Is the fairness of GNN models also highly vulnerable?” For example, in e-commerce, if attackers could exacerbate performance disparities between male and female user groups by attacking GNN-based recommendation models, they could ultimately cause the e-commerce platform to provoke dissatisfaction from specific user demographics and gradually lose its appeal among these users. Several studies [11, 13, 41] have explored the vulnerability of GNN fairness and proposed effective attack strategies. Unlike conventional attacks, these fairness attacks aim to undermine GNN fairness without excessively compromising its utility. However, all these works require altering connectivity between existing nodes, whose authority is typically limited in the real world, such as modifying the relationship between real users. In contrast, injecting fake nodes into the original graph is a more practical way to launch an attack without manipulating the existing graph [12, 32, 34], which is still under-explored in the field of GNN fairness attack. To address this gap, we aim to be the first to launch an attack on GNN fairness via node injection, examining their vulnerabilities under such a more realistic setting.
@@
Specifically, launching a node injection-based fairness attack on GNNs is non-trivial, whose challenges can be summarized as follows: RQ1: How to determine the node injection strategy? The node injection can be decomposed into two steps, including selecting appropriate target nodes and connecting the injected nodes with them, both will impact the effectiveness of the attack. RQ2: How to determine the features of the injected nodes after node injection? Like the real nodes, injected nodes will also participate in the message propagation process of GNNs, thereby affecting their neighbors and even the whole graph. Given the key role of massage propagation in GNN fairness [40, 45], proposing suitable strategies to determine the features of injected nodes is also important.
@@
To address these challenges, we propose a gray-box poisoning attack method namely Node Injection-based Fairness Attack (NIFA) during the GNN training phase. In detail, for the two steps in the first challenge, NIFA innovatively designs two corresponding principles. The first is the uncertainty-maximization principle, which asks to select real nodes with the highest model uncertainty as target nodes for injection. The idea is that nodes with higher uncertainty are typically more susceptible to attacks, thereby ensuring the attack’s effectiveness. After selecting target nodes, NIFA follows the second principle, the homophily-increase principle, to connect target nodes with injected nodes. This principle aims to deteriorate GNN fairness by enhancing message propagation within sensitive groups [22, 40]. For the second challenge, multiple novel objective functions are proposed after node injection to guide the optimization of the injected nodes’ features, which could further impact the victim GNN’s fairness from a feature perspective. In summary, our contributions are as follows:
@@ 2 Related work
Fairness on GNNs. Researchers have discovered various fairness issues of GNNs, which often lead to societal harms [22, 37] and performance deterioration [21, 33] in practical applications. Algorithmic fairness on GNNs can be categorized into two main types based on the definition: individual fairness [1, 6] and group fairness [5, 37, 46]. Individual fairness requires that similar individuals should receive similar treatment, while group fairness aims to protect specific disadvantaged groups [16]. In detail, many researchers have delved into studies focusing on fairness grounded in sensitive attributes. For instance, Dai et al. [5] reduce the identifiability of sensitive attributes in node embeddings through adversarial training to enhance fairness. FairVGNN [38] goes a step further by introducing a feature masking strategy to address the problem of sensitive information leakage during the feature propagation process in GNNs. Graphair [17] achieves fairness through an automated data augmentation method and FairSIN [40] designs a novel sensitive information neutralization method for fairness. Beyond fairness related to sensitive attributes, some researchers also direct attention to fairness related to graph structures, like DegFairGNN [21] and Ada-GNN [23]. In this work, we mainly focus on attacks on the group fairness of GNNs based on sensitive attributes.
@@
Attacks on GNNs. Finding out potential vulnerabilities thus improving the security of GNNs remains a pivotal concern in the field of trustworthy GNNs [42]. From the perspective of attackers, they aim to compromise the GNNs’ performance on graph data via manipulating graph structures [24, 32, 47], node attributes [48], or node labels [19]. Among these methods, node injection attacks, given the attackers’ limited authority to manipulate the connectivity between existing nodes, emerge as one of the most prevalent methods [32, 34, 43]. However, existing attacks, including node injection attacks, mainly focus on undermining GNN’s utility, with little attention to the vulnerability of GNN fairness. Different from attacks on GNN utility, fairness-targeted attacks aim to deteriorate the fairness without significantly compromising the accuracy. FA-GNN [11], FATE [13], and G-FairAttack [41] stand out as the few ones that we are aware of to explore attacks on GNN fairness. FA-GNN’s empirical findings suggest that adding edges with certain strategies can significantly compromise GNN fairness without affecting its performance [11]. FATE [13] formulates the fairness attacks as a bi-level optimization problem and proposes a meta-learning-based framework. G-FairAttack [41] designs a novel surrogate loss with utility constraints to launch the attacks in a non-gradient manner. Nevertheless, all these works require modifying the link structure between existing nodes, which may be prohibited in reality due to the lack of authority.
@@ 3 Preliminary
3.1 Notations
A graph is denoted as $\mathcal{G}=(\mathcal{V},\textbf{A},\textbf{X})$, where $\mathcal{V}$ is the node set, and $\textbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ represents the adjacency matrix. $\textbf{X}\in\mathbb{R}^{|\mathcal{V}|\times D}$ denotes the feature matrix, in which $D$ is the feature dimension. Under the settings of node classification, each node $v\in\mathcal{V}$ will be assigned with a label $y_{v}\in\mathcal{Y}$, and a GNN-based mapping function $f_{\theta}:\{\mathcal{V},\mathcal{G}\}\rightarrow\{1,2,...,|\mathcal{Y}|\}^{|\mathcal{V}|}$ with parameters $\theta$ is learned to leverage the graph signals for label prediction, where $\mathcal{Y}$ represents the true label set.
3.2 Fairness-related concepts
In alignment with prior works [5, 7, 17], we mainly focus on group fairness where each node will be assigned with a binary sensitive attribute $s\in\{0,1\}$, although our attack could also be generalized to the settings of multi-sensitive groups and we leave this as our future work. Based on the sensitive attributes, the nodes can be divided into two non-overlapped groups $\mathcal{V}=\{\mathcal{V}_{0},\mathcal{V}_{1}\}$, and we employ the following two kinds of fairness related definitions:
@@ 3.2 Fairness-related concepts
$\Delta_{SP}=\mathbb{E}|P(\hat{y}=y|s=0)-P(\hat{y}=y|s=1)|,$ (3)
$\Delta_{EO}=\mathbb{E}|P(\hat{y}=y|y=y,s=0)-P(\hat{y}=y|y=y,s=1)|.$ (4)
3.3 Problem definition
In this paper, our goal is to launch fairness-targeted attacks on GNN models through the application of node injection during the training phase, i.e. poisoning attack. Following the line of previous attacks on GNNs [32, 41], our attack is under the prevalent gray-box setting, where the attackers can obtain the graph $\mathcal{G}$ with node labels $\mathcal{Y}$, and the sensitive information $s$, but can not access the model architecture and parameters $\theta$. Detailed introduction to our attack settings is provided in Appendix B. Specifically, through injecting malicious node set $\mathcal{V}_{I}$ into the graph, the original graph $\mathcal{G}=(\mathcal{V},\textbf{A},\textbf{X})$ is poisoned as $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\textbf{A}^{\prime},\textbf{X}^{\prime})$, where
$\mathcal{V}^{\prime}=\mathcal{V}\cup\mathcal{V}_{I},\enspace\textbf{X}^{\prime}=\left[\begin{array}[]{cc}\textbf{X}\\ \textbf{X}_{I}\\ \end{array}\right],\textbf{X}_{I}\in{\mathbb{R}^{|\mathcal{V}_{I}|\times D}},$ (5)
$\textbf{A}^{\prime}=\left[\begin{array}[]{cc}\textbf{A}&\textbf{V}_{I}\\ \textbf{V}_{I}^{T}&\textbf{A}_{I}\\ \end{array}\right],\textbf{V}_{I}\in{\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}_{I}|}},\textbf{A}_{I}\in{\mathbb{R}^{|\mathcal{V}_{I}|\times|\mathcal{V}_{I}|}}.$ (6)
@@ 3.3 Problem definition
$\begin{array}[]{c@{}l}\mbox{s.t.}&\mathop{\arg\max}\limits_{\theta^{*}}\mathcal{M}(f_{\theta^{*}}(\mathcal{V},\mathcal{G}^{\prime})),\enspace\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\textbf{A}^{\prime},\textbf{X}^{\prime}),\enspace|\mathcal{V}_{I}|\leq b,\enspace deg(v)_{v\in\mathcal{V}_{I}}\leq d.\end{array}$
As a poisoning attack, the first constraint in Eq. (7) requires to train the victim model $f_{\theta^{*}}$ with parameters $\theta^{*}$ on the poisoned graph $\mathcal{G}^{\prime}$, so that the predictions of $f_{\theta^{*}}$ are as correct as possible before evaluating the attack performance. The following constraints in Eq. (7) make sure that the proposed attack is unnoticeable and deceptive to the defenders, i.e. the number of injected nodes is below a predefined budget $b$ [^1] Same as the prior work [32], we define the perturbation rate as the ratio of injected nodes to the labeled nodes in the original graph, i.e. $|\mathcal{V}_{I}|/|\mathcal{V}_{L}|$, where $\mathcal{V}_{L}$ denotes the labeled node set. and the degrees of injected nodes are constrained by a budget $d$. Our goal is to find a poisoned graph $\mathcal{G}^{\prime}$ to deteriorate the fairness of victim models $f_{\theta^{*}}$ as severely as possible, i.e. maximize the fairness metrics $\Delta_{SP}$ and $\Delta_{EO}$ introduced previously.